# The N3LO Twist-2 Matching of Linearly Polarized Gluon TMDs

### v2

@@ 2 Definition of linearly polarized gluon TMD PDFs and FFs
where $N(P)$ is a hadron state with momentum $P^{\mu}=(\bar{n}\cdot P)n^{\mu}/2=P^{+}n^{\mu}/2$, with $n^{\mu}=(1,0,0,1)$ and $\bar{n}^{\mu}=(1,0,0,-1)$, and $A_{n\perp}^{a,\mu}$ is the gauge invariant collinear gluon field with color index $a$ and Lorentz index $\mu$.
For sufficiently small $b_{\perp}$, the TMD PDFs in Eq. (1) admit operator product expansion onto the usual collinear PDFs,
@@
where $K^{\mu}$ is the total momentum of $|X_{n}\rangle$, and $d\text{PS}_{X_{n}}$ is the collinear phase space measure, and ${\rm\mathbb{P}}_{g\leftarrow i}^{\mu\nu}$ is the spin correlator of the color averaged gluon splitting amplitude
${\rm\mathbb{P}}_{g\leftarrow i}^{\mu\nu}\equiv{\rm\mathbb{Sp}}_{X_{n}g_{\mu}^{*}\leftarrow i}^{*}{\rm\mathbb{Sp}}_{X_{n}g_{\nu}^{*}\leftarrow i}\equiv-xP_{+}\langle i|{\cal A}_{n\perp}^{a,\mu}|X_{n}\rangle\langle X_{n}|{\cal A}_{n\perp}^{a,\nu}|i\rangle\,$, (12)
The integral in Eq. (11) is properly defined only with a rapidity cutoff implemented through the exponential rapidity regulator $e^{-b_{0}\tau\frac{P\cdot K}{P^{+}}}$ Li:2016axz ; Luo:2019hmp , which effectively restricts the total energies of the collinear radiations by a rapidity scale $\nu=1/\tau$.

### v3

@@ 2 Definition of linearly polarized gluon TMD PDFs and FFs
where $N(P)$ is a hadron state with momentum $P^{\mu}=(\bar{n}\cdot P)n^{\mu}/2=P^{+}n^{\mu}/2$, with $n^{\mu}=(1,0,0,1)$ and $\bar{n}^{\mu}=(1,0,0,-1)$, and $\mathcal{A}_{n\perp}^{a,\mu}$ is the gauge invariant collinear gluon field with color index $a$ and Lorentz index $\mu$. Equation (1) follows the normalization of the gluon fields in SCET label formalism. An equivalent, but more conventional definition of the gluon TMDs is expressed in terms of the QCD gluon field-strength tensor with a gauge link in the adjoint representation
${\cal B}_{g/N}^{{\rm bare},\mu\nu}(x,b_{\perp})=$ $\frac{1}{xP_{+}}\int\frac{db^{-}}{2\pi}e^{-ixb^{-}P^{+}}\langle N(P)|G^{+\mu}(0,b^{-},b_{\perp})\mathcal{W}(0,b^{-},b_{\perp};0)G^{+\nu}(0)|N(P)\rangle\,.$ (2)
For sufficiently small $b_{\perp}$, the TMD PDFs in Eq. (1) admit operator product expansion onto the usual collinear PDFs deFlorian:2001zd ; Catani:2011kr ; Catani:2012qa ; Gutierrez-Reyes:2017glx ; Gehrmann:2012ze ; Gehrmann:2014yya ; Echevarria:2016scs ; Gutierrez-Reyes:2019rug ; Luo:2019hmp ; Luo:2019bmw ; Catani:2022sgr ; Ebert:2020qef ; Ebert:2020yqt ; Luo:2020epw
@@
where $K^{\mu}$ is the total momentum of $|X_{n}\rangle$, and $d\text{PS}_{X_{n}}$ is the collinear phase space measure, and ${\rm\mathbb{P}}_{g\leftarrow i}^{\mu\nu}$ is the spin correlator of the color averaged gluon splitting amplitude Bern:1993qk ; Bern:1994zx ; Bern:1995ix ; Kosower:1999xi ; Bern:1999ry ; Catani:1999ss ; Catani:2003vu ; Catani:2011st ; Bern:1998sc ; Kosower:1999rx ; Sborlini:2013jba ; DelDuca:1999iql ; Birthwright:2005ak ; Birthwright:2005vi ; DelDuca:2020vst ; DelDuca:2019ggv ; Badger:2015cxa ; Czakon:2022fqi ; Bern:2004cz ; Badger:2004uk ; Duhr:2014nda
${\rm\mathbb{P}}_{g\leftarrow i}^{\mu\nu}\equiv{\rm\mathbb{Sp}}_{X_{n}g_{\mu}^{*}\leftarrow i}^{*}{\rm\mathbb{Sp}}_{X_{n}g_{\nu}^{*}\leftarrow i}\equiv-xP_{+}\langle i|{\cal A}_{n\perp}^{a,\mu}|X_{n}\rangle\langle X_{n}|{\cal A}_{n\perp}^{a,\nu}|i\rangle\,.$ (13)
The integral in Eq. (12) requires a rapidity cutoff to be well defined Collins:1984kg ; Ji:2004wu ; Collins:2011zzd ; Becher:2010tm ; Becher:2011dz ; Chiu:2012ir ; Chiu:2009yx ; Echevarria:2015byo ; Li:2016axz . In this work, we implement such a cutoff via the exponential rapidity regulator $e^{-b_{0}\tau\frac{P\cdot K}{P^{+}}}$ Li:2016axz ; Luo:2019hmp , which effectively restricts the total energies of the collinear radiations by a rapidity scale $\nu=1/\tau$.